can you please help me with. find the value of the function.
find f(x-1) when f(x)=4x^2+3x+5

Answer by stanbon(75887)

find f(x-1) when f(x)=4x^2+3x+5
-------------
f(x-1) = 4(x-1)^2+3(x-1)+5
-----
= 4(x^2-2x+1)+3x-3+5
---
= 4x^2-8x+4+3x+2
----
= 4x^2-5x+6
================
cheers,
Stan H.